OpenStudy (anonymous):

A square has a side of 6.25 feet. What is the area in square feet? Round to the correct significant digits.

jigglypuff314 (jigglypuff314):

6.25 has three sig figs area of square = length time length so = 6.25 * 6.25 then round to three sig figs

OpenStudy (anonymous):

A square has all four sides equal in length to each other. Area = side x side Area = 6.25 * 6.25 = 39.0625 convert to 2 sig dig Area = 39.06 square feet.

OpenStudy (anonymous):

what does it mean by round to the significant digits

OpenStudy (anonymous):

round to the nearest tenth

OpenStudy (anonymous):

40

jigglypuff314 (jigglypuff314):

kinda. 6.25 has three sig figs ^ ^^ three digits total therefore round to a number with 3 digits total so 6.25 times 6.25 = 39.0625 -> 39.1

OpenStudy (anonymous):

i get it now thanks!

jigglypuff314 (jigglypuff314):

you're welcome :)